OpenStudy (anonymous):

√X + Y = 7 X + √Y = 11 find x and y!

OpenStudy (anonymous):

x-y = 7 y = 2

OpenStudy (anonymous):

Haha! Nop! Wrong! (:

OpenStudy (anonymous):

are both positive?

OpenStudy (anonymous):

there are 2 solutions, one of them would be x=8.51253 and y=6.1875

OpenStudy (anonymous):

You want the solutions to the two equations?

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (akashdeepdeb):

The solutions are very easy to find with hit and trial. But I am sure he needs the method. Using a calculator is cheating here!

OpenStudy (anonymous):

The solutions are easy its X=9 and Y=4! :D What is the method here? @greenlegodude57 @AkashdeepDeb

OpenStudy (anonymous):

From the first equation, \[\sqrt x+y=7~~\iff~~\sqrt x=7-y~~\Rightarrow~~x=(7-y)^2\] Substitute into the second equation: \[(7-y)^2+\sqrt y=11\] Expand and simplify: \[\begin{align*}49-14y+y^2+\sqrt y&=11\\ \sqrt y&=-y^2+14y-38\\ y&=\left(y^2-14y+38\right)^2\\ 0&=y^4-28y^3+272y^2-1065y+1444 \end{align*}\] \(y=4\) is a root, but how to get that manually? Maybe try the rational root theorem, or guess and check, noting that \(x,y>0\) must hold.